# StefaLand: An Efficient Geoscience Foundation Model That Improves Dynamic Land–Surface Predictions

### v1

@@ 2.5 Foundation model comparisons
For completeness, we also evaluated two existing Earth-Observation-oriented foundation models, TerraMind and PrithviWxC Jakubik et al. (2025), Hsu et al. (2024), using the same fine-tuning heads and training protocols as StefaLand. TerraMind was tested on streamflow and soil moisture, and PrithviWxC on soil moisture. Inputs were harmonized to each model’s modality. It is noteworthy that either Terramind or PrithviWxC requires orders of magnitude larger storage space for the input data than StefaLand, as well as computational cost for finetuning. Due to intensive data requirements for PrithviWxC on surface-level variables mostly likely to demonstrate relevance to land surface interactions, along with all 14 static variables were used. The full atmospheric variables at differing pressure levels were excluded.
@@ 3 Experiments
For all experiments described below, forcing data and static features remained consistent across all model configurations, with hyperparameters held constant within each test type except for fundamental architectural differences inherent to transformer models. Complete details of hyperparameters, forcings, and static features are provided in the Appendix C.
In addition to the neural and process-based baselines reported below, we trained plain linear regression models for every task using the same features (and appropriate standardization). These consistently underperformed all learning baselines by large margins, so we omit them from tables for readability; full metrics are provided in Appendix E.
@@ 3.1 CAMELS Streamflow Prediction
To compare the scheme’s spatial generalization capability on a well benchmarked dataset, we follow the experimental setup in (Feng et al., 2021), testing prediction in ungauged basins (PUB) and ungauged regions (PUR). These correspond to randomized spatial K-fold and regional-specific K-fold regimes, reflecting real-world challenges in generalizing to nearby unmonitored basins and to data-scarce regions, respectively. We use the CAMELS dataset (Addor et al., 2017, Newman et al., 2014), restricted to the 531-basin subset with clear watershed boundaries (Newman et al., 2017). Basins were divided into 10 random groups for PUB and 7 contiguous regions for PUR, employing leave-one-out in both cases. To avoid leakage, all CAMELS-overlapping stations were removed during pretraining for PUB tests, and entire regions were excluded for PUR tests. We evaluated both direct prediction models and neural networks used to parameterize process-based models (see Methods). Direct models included StefaLand without finetuning (StefaLand-Transformer), pretrained variants with and without residual connections (resConn / noResConn), a from-scratch ablation trained only on task data, and a version using TerraMind as encoder. For process-based models, we tested LSTM HBV1.1, StefaLand-resConn HBV1.1, and StefaLand-noResConn HBV1.1.
@@
RMSE $\downarrow$ pbRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$ RMSE $\downarrow$ pbRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$
Direct Prediction Models
@@
StefaLand Ablation - resConn 1.355 1.332 0.801 0.661 1.516 1.378 0.771 0.560
Physics Based Models
LSTM - HBV1.1 1.325 1.298 0.857 0.672 1.561 1.521 0.746 0.578
The original multi-basin LSTM studies on CAMELS defined the modern supervised learning benchmark (Kratzert et al., 2019), and subsequent large-scale comparisons confirmed that vanilla Transformers generally fail to outperform LSTMs on rainfall–runoff prediction (Liu et al., 2024, Table 1 therein). This is again reinforced in our results, where StefaLand-Transformer underperforms LSTM. The Nash-Sutcliffe model efficiency coefficient (NSE) values reported here for LSTM are highly similar to those reported in the domain literature Feng et al. (2021). Broader evaluations likewise continue to rank LSTM-family models among the best-performing approaches (Lees et al., 2021), and even Google’s global flood-forecasting system adopts an encoder–decoder LSTM backbone (Nearing et al., 2024). We therefore use LSTM-SL as the supervised SOTA comparator, complemented by a process-informed LSTM–HBV1.1 variant, and additionally test TerraMind to control for the role of EO-centric pretraining.
Against these strong baselines, the foundation model pretraining clearly provides a rare boost to generalization capability compared to supervised learning models (LSTM-SL and StefaLand-Transformer) across both PUB and PUR scenarios (Table 1). StefaLand-resConn’s RMSE is almost $20\%$ lower than that of supervised LSTM in the PUB and $17\%$ lower in the harder PUR test, while the correlation are noticebly higher. The value of pretraining is confirmed by StefaLand-Ablation-resConn, which appears to have only minor advantage compared to LSTM. The residual-connection architecture enables effective integration of transformer features with temporal dynamicsh, showing a modest benefit compared to StefaLand-no resConn, but both are clearly stronger than either LSTM or StefaLand-Transformer. Finally, Terramind-resConn has quite comparable metrics to LSTM. It appears the pretrained Terramind offers no generalization value.
The foundation model showed its versatility in supporting the parameterization of hybrid models and improving their generalization. We observe the same patterns — the pretrained StefaLand is helpful and resConn is a valuable construct. The physics-based experiments highlight that residual connections can combine pretrained representations with process-based constraints, producing more reliable spatial generalization than even the strongest LSTM configurations. We note that the hybrid model in fact showed a slightly better NSE than the direction prediction models for PUR for either LSTM-HBV1.1 or StefaLand-resConn-HBV1.1, because the physics-based equations serve as additional constraints to mitigate overfitting.
@@ 3.2 Global Streamflow
We designed a global-scale runoff prediction experiment to assess the robustness and generalization ability worldwide. We filtered global basin datasets based on data completeness and retained 3,434 basins for the global experiment. To manage computational costs efficiently, we employed random hold-out sampling combined with three-fold cross-validation. Additionally, we implemented a regionally hold-out continental scenario (RH-C), excluding all basins from North America, South America, and Europe, respectively, from the training set. The model trained from remaining continents was then evaluated on the excluded continent, simulating the realistic scenario of deploying the model to unseen geographical domains.
@@
StefaLand - Transformer 0.749 0.751 0.843 1.075 1.048 0.697
Terramind + finetuning 1.156 1.111 0.580 1.234 1.158 0.670
@@
Results indicate that StefaLand-Transformer outperformed both LSTM-SL and the fine-tuned vision-based Transformer model (TerraMind) across all evaluation metrics. Specifically, in the RH scenario, StefaLand achieved an RMSE of 0.749, approximately 14% lower than LSTM (0.87), with Corr improving to 0.843. Notably, StefaLand’s unbiased RMSE closely matched its RMSE, indicating errors were primarily random fluctuations rather than systematic bias. In contrast, TerraMind exhibited a higher RMSE of 1.156 and a low Corr of 0.580. We hypothesize that this is because the pretraining satellite image data for the vision foundation network TerraMind, did not have high relevance to hydrologic predictions.
Under the more challenging RH-C scenario, all models experienced increased errors, but StefaLand continued to demonstrate higher performance. Its RMSE was 1.075, approximately 14.1% lower than LSTM’s 1.253, while also maintaining the highest Corr (0.697). Moreover, StefaLand’s ubRMSE remained close to its RMSE, confirming its robust ability to correct regional-scale biases even under extreme out-of-domain conditions. Conversely, LSTM and TerraMind exhibited large gaps between ubRMSE and RMSE, highlighting their challenging in producing hydrologically-relevant features.
@@ 3.3 Global Soil Moisture
We evaluated finetuning StefaLand for soil moisture predictions following the experimental design from Liu et al. (2023a), using data from the International Soil Moisture Network (ISMN) Dorigo et al. (2011; 2013). Even though there is a globally covering satellite-based product for soil moisture, the data quality can hardly match that of in-situ moisture sensors. Thus the ability to generalize in-situ data will still be valuable. ISMN consists of 1,316 ground-based measurement stations. We randomly partitioned sites into five groups for spatial cross-validation in the random holdout test, training on four groups and testing on the fifth, rotating through all groups to obtain comprehensive spatial generalization performance. For the regional holdout test, we specifically evaluated model performance on Europe as an ungauged region, training on all other continents while excluding European sites entirely to assess cross-continental transferability, removing 129 sites for testing. We tested the following model configurations: StefaLand-Transformer (no finetuning, direct prediction), StefaLand with and without residual connections (resConn / noResConn), an ablation trained from scratch without pretraining, a baseline LSTM, and a version using IBM’s TerraMind encoder with resConn.
@@
PrithviWcX - resConn 0.081 0.060 0.703 0.103 0.079 0.523
Ablation StefaLand - resConn 0.074 0.058 0.749 0.108 0.064 0.528
@@
LSTM again serves here as the established state-of-the-art baseline. Comprehensive reviews have shown that LSTM variants consistently provide the most stable and effective performance among supervised learning methods for soil moisture prediction due to their temporal modeling strengths (Wang et al., 2024). On the ISMN-linked global dataset in particular, multitask LSTM models are widely adopted and competitive with newer architectures (Liu et al., 2023b). Given these findings, we benchmark against LSTM-SL as the SOTA/near-SOTA supervised model, and include transformer-only and vision-foundation encoders to disentangle temporal modeling from representation learning.
Against these strong baselines, the soil moisture experiments confirm the superiority of the StefaLand-resConn architecture (Table 3). StefaLand-resConn achieves the best performance, with RMSE of 0.068 and correlation of 0.783 for random holdout, and maintains strong performance even in cross-continental testing on Europe (RMSE = 0.090, Corr = 0.638). The direct StefaLand model performs poorly since transformers alone lack an effective mechanism for temporal dependencies, while the resConn architecture compensates for this limitation. Similar to the streamflow results, TerraMind and PrithviWcX both showed no performance benefit compared to LSTM despite architectural adaptations, confirming both the challenge of repurposing vision foundation models across domains and the importance of pretraining with geoscience-relevant variables. Notably, the regional holdout on Europe demonstrates StefaLand-resConn’s superior spatial generalization capabilities, achieving a 25% correlation improvement over the LSTM baseline in this highly challenging extrapolation scenario.
We want to emphasize that TerraMind and PrithviWxC were not designed for the hydrologic and land–surface prediction tasks studied here. They excel in their intended domains of EO imagery and atmospheric variables, but are out-of-domain for dynamic land–surface and hydrologic modeling. We include them only to explore whether such models transfer any useful signal in our setting. Their weaker results should not be interpreted as evidence against their capability in other applications, but rather as a reflection of the mismatch between their pretraining objectives and the land–surface tasks we target. Nevertheless, we believe these benchmarks are helpful for clarifying their respective strengths, since the AI community may not be familiar with these datasets and models that have been claimed "any-to-any generative foundation model for Earth observation".
@@ 3.4 Soil Property Prediction
There are different soil datasets, each collected with different protocols and data processing techniques, resulting in significant discrepancies between datasets. In this test, we finetuned StefaLand to predict in-situ soil profile data from another dataset (ISRIC). This application can produce a seamless dataset that is consistent with a set of in-situ data, improving data availability and addressing systematic biases. In addition, it helps us understand the noise associate with each dataset. StefaLand’s pretraining soils dataset is HWSD, which has some overlap but also quite extensive differences from ISRIC, which is larger and potentially noisier. We finetuned StefaLand to predict one soil texture property (e.g., clay percentage) in ISRIC while masking the corresponding complementary attributes (e.g., sand, silt) from the same profile to avoid information leakage. As a result, we are probing how easy it is to infer soil properties using other attributes such as climate, terrain and land cover. We compared StefaLand with and without pretraining against a supervised random forest baseline, which is prevalently used in the present geoscientific literature.
@@
StefaLand - Transformer Clay 0.20 0.04
StefaLand - Finetune Clay 0.51 0.26
Random Forest Clay 0.05 0.00
StefaLand - Transformer Sand 0.25 0.06
StefaLand - Finetune Sand 0.70 0.50
Random Forest Sand 0.26 0.07
StefaLand - Transformer Silt 0.41 0.17
StefaLand - Finetune Silt 0.63 0.40
Random Forest Silt 0.84 0.70
StefaLand with finetuning achieved markedly higher predictive power, especially for sand fraction ($R^{2}=0.50$), compared to both direct training and the random forest baseline. These results highlight StefaLand’s ability to reconcile disparate soil datasets and improve the utility of noisy in-situ observations. At the same time, the weaker performance on silt suggests limits of transfer: silt fractions are often residually defined (sand and clay measured directly, silt inferred as the remainder), and thus more sensitive to error propagation across datasets. This makes them inherently harder to learn, as any noise or systematic difference between HWSD and ISRIC is amplified. The fact that the random forest attains higher accuracy on silt may also indicate that simpler models can more directly exploit the compositional closure relationship, whereas StefaLand relies more on cross-variable and environmental cues, which provide less direct signal for silt.

### v2

@@ 2.4 Foundation model comparisons
For completeness, we also evaluated two existing Earth-Observation-oriented foundation models, TerraMind and PrithviWxC (Jakubik et al., 2025, Hsu et al., 2024), using the same fine-tuning heads and training protocols as StefaLand. We coupled them together with our residual connection architecture and finetuned the added units while keeping the foundation models’ weights frozen. Because both TerraMind and PrithviWxC require orders of magnitude larger storage space, data transfer, and fine-tuning costs than StefaLand, we chose comparison tasks selectively, focusing on cases where their pretraining data was potentially most relevant. Accordingly, we compared both models on soil moisture, while streamflow was included only in a more exploratory capacity (only TerraMind). For a rough comparison, StefaLand, TerraMind and PrithviWxC used respectively $\approx 2$, $\approx 11$ and $\approx 27$ TB data during pretraining. Due to the intensive data requirements for PrithviWxC, only surface-level variables likely relevant to land surface interactions, along with all 14 static variables, were used. The full atmospheric variables at differing pressure levels were excluded.
@@ 3 Experiments
We tested the value of foundation model pretraining on 5 datasets and 6 experiments, including, streamflow on the CAMELS dataset on USA, CAMELS streamflow prediction with hybrid model, global streamflow, global in-situ soil moisture, global soil properties, and landslide susceptibility in Oregon, USA. For all experiments, hyperparameters were tuned with Ray Tune and kept consistent across model configurations within each experimental case (e.g., CAMELS streamflow, soil moisture, etc). Because we compare spatial generalization, we used temporal validation splits for hyperparameter optimization. Complete details of hyperparameters, forcings, and static features for all four experiments are in Appendix C.
Model variants used.
Unless noted otherwise, pretrained encoders are frozen and only task heads are trained. We evaluate: (i) LSTM-SL: supervised LSTM baseline; (ii) StefaLand-direct: the StefaLand encoder together with its original decoder trained directly on the task data, without pretraining; (iii) StefaLand-resConn: pretrained encoder with residual pathway + LSTM decoder, our proposed architecture; (iv) StefaLand-noResConn: pretrained encoder with a simple adapter (no residual/LSTM); (v) StefaLand-scratch: same resConn head but StefaLand initialized randomly (no pretraining) and unfrozen (ablation test); (vi) TerraMind-resConn and PrithviWxC-resConn: EO/atmospheric foundation encoders frozen with the same residual head.
@@ 3.1 CAMELS Streamflow Prediction
To compare spatial generalization on a well benchmarked dataset, we follow (Feng et al., 2021), testing prediction in ungauged basins (PUB) and ungauged regions (PUR). These correspond to randomized spatial K-fold and regional-specific K-fold regimes, respectively. We use CAMELS (Addor et al., 2017, Newman et al., 2014), restricted to the 531-basin subset with clear watershed boundaries (Newman et al., 2017). Basins were divided into 10 random groups for PUB and 7 contiguous regions for PUR, employing leave-one-out in both cases. To avoid leakage, all CAMELS-overlapping stations were removed during pretraining for PUB, and entire regions were excluded for PUR.
@@
RMSE $\downarrow$ µbRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$ RMSE $\downarrow$ µbRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$
@@
The foundation model pretraining clearly provides a rare boost to generalization capability compared to supervised learning models (LSTM-SL and StefaLand-direct) across both PUB and PUR (Table 1). StefaLand-resConn’s RMSE is almost $20\%$ lower than that of supervised LSTM in PUB and $17\%$ lower in PUR, while the correlation are noticeably higher. The value of pretraining is confirmed by StefaLand-scratch, which appears to have only minor advantage compared to LSTM. The residual-connection architecture enables effective integration of transformer features with temporal dynamics, showing a modest benefit compared to StefaLand-no resConn, but both are clearly stronger than either LSTM or StefaLand-direct. Finally, TerraMind-resConn has quite comparable metrics to LSTM; the pretrained TerraMind offers no extra generalization value in this case.
We ran additional experiments that hybridize StefaLand with the HBV1.1 physics backbone on the same PUB/PUR splits, testing its ability to parameterize physics-based models. These hybrids achieved up to a 13% RMSE reduction and a 10% correlation gain compared to the LSTM–HBV1.1 baseline, demonstrating that pretraining strengthens hybridization with process models. By constraining predictions with physics while leveraging StefaLand features, these hybrids further improve upon the general results above and highlight the versatility of the approach. Full results are provided in Appendix B, Table B1.
On a related note, the supervised LSTM is not an easy benchmark to surpass. The original multi-basin LSTM and subsequent large-scale comparisons (Kratzert et al., 2019, Feng et al., 2021) showed that vanilla Transformers generally fail to outperform LSTMs on rainfall–runoff prediction (Liu et al., 2024, Table 1 therein). The LSTM NSE values reported here are very similar to those in the domain literature (Feng et al., 2021). Broader evaluations likewise continue to rank LSTM-family models among the best-performing approaches (Lees et al., 2021), and even Google’s global flood-forecasting system adopts an encoder–decoder LSTM backbone (Nearing et al., 2024).
@@ 3.2 Global Streamflow
We designed a global-scale runoff prediction experiment to assess robustness and generalization worldwide. We filtered global basin datasets based on data completeness and retained 3,434 basins. To manage computational cost, we employed random hold-out sampling with three-fold cross-validation. Additionally, we implemented a regionally hold-out continental scenario (RH-C), excluding all basins from North America, South America, and Europe, respectively, from training, then evaluating on the excluded continent.
@@
StefaLand - direct 0.749 0.751 0.843 1.075 1.048 0.697
TerraMind - resConn 1.156 1.111 0.580 1.234 1.158 0.670
@@
Results indicate that StefaLand-direct outperformed both LSTM-SL and the fine-tuned vision-based Transformer model (TerraMind) across all metrics for either random or regional holdout (Table 2). Specifically, in the random holdout scenario, StefaLand achieved an RMSE of 0.749, approximately 14% lower than LSTM (0.87), with Corr improving to 0.843. Notably, StefaLand’s unbiased RMSE closely matched its RMSE, indicating errors were primarily random fluctuations rather than systematic bias. In contrast, TerraMind exhibited a higher RMSE of 1.156 and a low Corr of 0.580. We hypothesize this is because the pretraining satellite image data for the vision foundation network TerraMind did not have high relevance to hydrologic predictions like streamflow. We note that using TerraMind here was exploratory, as it was not specifically designed or pretrained for this use case.
Under the more challenging RH-C scenario, all models exhibited larger errors, but StefaLand continued to perform best. Its RMSE was 1.075, approximately 14.1% lower than LSTM’s 1.253, while also maintaining the highest Corr (0.697). Moreover, StefaLand’s ubRMSE remained close to its RMSE, confirming robust correction of regional-scale biases even under extreme out-of-domain conditions. Conversely, LSTM and TerraMind showed larger gaps between ubRMSE and RMSE, highlighting challenges in producing hydrologically-relevant features.
@@ 3.3 Global Soil Moisture
We evaluated finetuning StefaLand for soil moisture predictions following Liu et al. (2023a), using ISMN (Dorigo et al., 2011; 2013a). Even though there is a globally covering satellite-based product for soil moisture, the data quality can hardly match that of in-situ moisture sensors; thus the ability to generalize in-situ data is valuable. ISMN consists of 1,316 ground-based stations. We performed five-fold spatial cross-validation for random holdout and a regional holdout on Europe, training on all other continents while excluding European sites (129) for testing. We tested StefaLand-direct, StefaLand with and without residual connections (resConn / noResConn), a from-scratch ablation, an LSTM baseline, and a version using IBM’s TerraMind encoder with resConn. LSTM again serves as the established state-of-the-art baseline (Wang et al., 2024, Liu et al., 2023b).
@@
PrithviWxC - resConn 0.081 0.060 0.703 0.103 0.079 0.523
@@
Against these strong baselines, the soil moisture experiments confirm the superiority of the StefaLand-resConn architecture (Table 3). It achieves the best performance, with RMSE of 0.068 and correlation of 0.783 for random holdout, and maintains strong performance even in cross-continental testing on Europe (RMSE = 0.090, Corr = 0.638). The direct StefaLand model performs poorly because it lacks an effective mechanism to capture task-specific temporal dependencies. Similar to the streamflow cases above, TerraMind and PrithviWxC showed no performance benefit compared to LSTM, underscoring the challenge of repurposing vision foundation models across domains and the importance of pretraining with geoscience-relevant variables. The regional holdout on Europe further demonstrates StefaLand-resConn’s superior spatial generalization, achieving a 25% correlation improvement over the LSTM baseline in this difficult extrapolation scenario.
We want to emphasize that TerraMind and PrithviWxC were not designed for the hydrologic and land–surface prediction tasks studied here. They excel in their intended domains of Earth observation imagery and atmospheric variables, but are out-of-domain for dynamic land–surface and hydrologic modeling. We include them only to explore whether such models transfer any useful signal in our setting. These results should not be interpreted as evidence against their capability in other applications, but rather as a reflection of the mismatch between their pretraining objectives and the land–surface tasks we target. Nevertheless, we believe these benchmarks are helpful for clarifying their respective strengths, since the general AI community may not be familiar with these datasets and models that have been marketed as ‘any-to-any’ generative EO foundation models. Ultimately we view StefaLand as complementary to satellite-image foundation models, targeting a different but relevant data regime. We also could not identify other foundation models for land surface.
@@ 3.4 Soil Property Prediction
There are different soil datasets, each collected with different protocols and data processing techniques, resulting in significant discrepancies. In this test, we finetuned StefaLand to predict in-situ soil profile data from another dataset (ISRIC). This application can produce a seamless dataset that is consistent with a set of in-situ data, improving data availability and addressing systematic biases. In addition, it helps us understand the noise associated with each dataset. StefaLand’s pretraining soils dataset is HWSD, which has some overlap but also extensive differences from ISRIC, which is larger and potentially noisier. We finetuned StefaLand to predict one soil texture property (e.g., clay percentage) in ISRIC while masking the corresponding complementary attribute (e.g., sand) from the same profile to avoid information leakage, probing how easy it is to infer soil properties using other attributes such as climate, terrain and land cover. We compared StefaLand with and without pretraining against a supervised random forest baseline. Train-test splits can be found here C.2
@@
StefaLand - direct Clay 0.197 0.038
StefaLand - finetune Clay 0.509 0.259
Random Forest Clay 0.456 0.207
Linear Regression Clay 0.138 0.019
StefaLand - direct Sand 0.253 0.064
StefaLand - finetune Sand 0.704 0.495
Random Forest Sand 0.585 0.342
Linear Regression Sand 0.347 0.120
StefaLand with finetuning achieved markedly higher predictive power for both clay and sand fractions, substantially outperforming direct training, random forest, and linear regression baselines. The other models struggle in this task, with random forest (RF) scoring noticeably lower. This task is not dynamical prediction, but a test of the models ability to build deep representations of the landscape that can be adapted to infer certain attributes using cross-variable group connections. StefaLand-finetune’s advantage against random forest suggests there are indeed deep representations uncovered by pretraining but not exploited by random forest. Finetuning StefaLand is also apparently superior to training from scratch as supervised learning (StefaLand - direct), again highlighting the value of pretraining. These results highlight StefaLand’s ability to reconcile disparate attribute datasets and improve the utility of noisy in-situ observations.
3.5 Landslide Susceptibility Prediction
Landslide is a geohazard that kill thousands each year. We next evaluated StefaLand for landslide susceptibility prediction using the SLIDO dataset from the State of Oregon, which provides detailed landslide occurrence records. Following Liu et al. (2025), this is a binary classification task indicating the presence or absence of landslides in a 30m by 30m patch. We finetuned StefaLand by extracting frozen hidden features and concatenating them with a 2D CNN, then retrained the CNN classifier to assess StefaLand’s ability to provide generalizable geoscience features.
Table 5: Landslide susceptibility prediction results on the Oregon SLIDO dataset.
Model Accuracy $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ ROC AUC $\uparrow$
Logistic Regression 0.742 0.707 0.792 0.47 0.819
Random Forest 0.751 0.703 0.834 0.763 0.839
CNN2D 0.778 0.760 0.787 0.773 0.863
StefaLand + CNN2D 0.799 0.791 0.793 0.792 0.875
Note: All baseline results (Logistic Regression, Random Forest, CNN2D) are taken from previously published 30m-resolution experiments in Liu et al. (2025), except for StefaLand + CNN2D, which represents our proposed method.
Results show that StefaLand’s pretrained features improved the CNN’s generalization, yielding modest gains across all metrics except Recall. This is a particularly difficult baseline to improve so even modest gains are rare. ROC AUC increased from 0.863 to 0.875, and precision rose from 0.760 to 0.791, reflecting fewer false positives. While random forest finds a high Recall, this is a particular realization with a large tradeoff on Precision. StefaLand produces overall well-rounded predictions, with both better Precision and Recall than CNN2D.